OpenStudy (anonymous):

i forget how to find what A can you help ? Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary. a = ?, b = 5, c = 8 A. 6.0 B. 6.1 C. 6.2 D. 6.3

OpenStudy (saifoo.khan):

do you have a particular diagram?

OpenStudy (saifoo.khan):

which is the longest side?

OpenStudy (anonymous):

\[a=\sqrt{64 - 25}\]

OpenStudy (anonymous):

You subtract b squared from c square so: 64-25 which equals = 39. On a calculator the square root of 39 is 6.2

OpenStudy (anonymous):

the longest side is c and no there is no diagram

OpenStudy (saifoo.khan):

\[8^2=5^2+a^2\]

OpenStudy (anonymous):

thank you great help